TheJag18:

A $225,000 adjustable-rate mortgage is expected to have the following payments: Year Interest Rate Monthly Payment 1–5 4% $1,074.18 6–15 6% $1,311.20 16–25 8% $1,484.91 26–30 10% $1,555.99 A fixed-rate mortgage in the same amount is offered for 30 years and 7% interest. Part A: What is the total cost of the adjustable-rate mortgage? State which technology or method you used. Show or explain all steps. (3 points) Part B: What is the total cost of the fixed-rate mortgage? State which technology or method you used. Show or explain all steps. (3 points) Part C: Using your values from parts A and B compare the advantages and disadvantages of the two loan types. (4 points)

TheJag18:

This is my answer

Aratox:

Part A: To calculate the total cost of the adjustable-rate mortgage, we need to add up all the monthly payments for each period. For the first 5 years: Total cost = $1,074.18 * 12 months/year * 5 years = $64,450.80 For the next 10 years: Total cost = $1,311.20 * 12 months/year * 10 years = $157,344.00 For the next 10 years: Total cost = $1,484.91 * 12 months/year * 10 years = $178,189.20 For the final 5 years: Total cost = $1,555.99 * 12 months/year * 5 years = $93,359.40 Adding all these costs together: Total cost of the adjustable-rate mortgage = $64,450.80 + $157,344.00 + $178,189.20 + $93,359.40 = $493,343.40 I used simple multiplication to calculate the total cost of the adjustable-rate mortgage. Part B: To calculate the total cost of the fixed-rate mortgage, we use the formula for a fixed-rate mortgage: Total cost = Monthly payment * 12 months/year * 30 years Total cost = $1,622.27 * 12 months/year * 30 years = $584,017.20 I used simple multiplication to calculate the total cost of the fixed-rate mortgage. Part C: Advantages of adjustable-rate mortgage: - Lower initial monthly payments, making it more affordable in the short term - Potential for lower interest rates and payments in the future if interest rates decrease Disadvantages of adjustable-rate mortgage: - Payments can increase significantly if interest rates rise - Uncertainty about future payments can make budgeting difficult Advantages of fixed-rate mortgage: - Predictable monthly payments for the entire term of the loan - Protection against rising interest rates Disadvantages of fixed-rate mortgage: - Higher initial interest rate compared to adjustable-rate mortgage - Higher total cost over the life of the loan In this case, the adjustable-rate mortgage has a lower total cost compared to the fixed-rate mortgage. However, it also comes with the risk of increasing payments if interest rates rise. The fixed-rate mortgage offers predictability and protection against rising interest rates, but at the cost of a higher total repayment amount. Ultimately, the choice between the two loan types will depend on individual preferences and risk tolerance.